# Entanglement entropy of two disjoint intervals and spin structures in interacting chains in and out of equilibrium

### 15 Dec 2023 v1

@@ 1 Introduction
An important aspect that distinguishes the entropy of disjoints blocks from that of an interval is its sensitivity to duality mappings, in that such transformations generally preserve the locality of only a subset of operators (including the Hamiltonian density), breaking it instead into semilocality for the rest of them. This affects the meaning itself of subsystem, especially when it is not simply connected. A consequence of it on the behaviour of the entanglement entropies was firstly observed in the ground state of XY spin chains Igloi2010 ; Fagotti2010disjoint . At the critical point the tripartite information has a nontrivial scaling limit despite vanishing in the corresponding free-fermion model that emerges after a Jordan-Wigner transformation Casini2005 ; Casini2009reduced_density ; Casini2009Entanglement .
@@
Since the first step towards an analytical investigation into the entanglement entropies of spin blocks is often a mapping to quasiparticles, it is important to understand how such mappings affect the entropies and how to ultimately fix the lack of equivalence. In that respect, Ref. Fagotti2010disjoint proposed an expansion of the tripartite information based on the decomposition of the reduced density matrix into four pseudo reduced density matrices (pseudo-RMDs) that unfold the spin-flip symmetry. On the other hand, the critical spin-chain models studied in Ref. Fagotti2010disjoint admit a quantum field theory description of the low-energy and large-distance limit (the Ising model is described by a massless Majorana field theory, while the XX model by a Dirac one); as mentioned above, the Rényi entropy $S_{\alpha}$ of two disjoint intervals can then be written in terms of partition functions on particular Riemann surfaces. Ref. Coser2016Spin examined what the lattice procedure of Ref. Fagotti2010disjoint translates into in the quantum field theory description and established a one-to-one correspondence between the partition functions with given spin structures and the moments of the pseudo-RDMs introduced in Ref. Fagotti2010disjoint . It was argued, in particular, that the entropy of disjoint blocks of spins is obtained by summing over all spin structures of the fermion model on the underlying Riemann surface, i.e., over all combinations of antiperiodic, also called Neveu-Schwarz (NS), and periodic, also called Ramond (R), boundary conditions on the cycles of the surface. The entropy of Jordan-Wigner fermions, instead, corresponds to antiperiodic boundary conditions over all cycles (see also Ref. Camargo2021 ). As discussed in Ref. Headrick2010 in detail, indeed, the discrepancy between the entanglement entropy of two disjoint intervals for the free Dirac fermion Casini2005 ; Casini2009reduced_density ; Casini2009Entanglement and for the free boson compactified on a circle with corresponding dual radius Calabrese2009Entanglement1 , $R=1$, can be traced back to the spin structure. The result for the boson matches, instead, that for the modular invariant Dirac fermion Calabrese2011Entanglement ; Coser2016Spin , which is obtained by summing over all boundary conditions. We mention that similar problems related to Bose-Fermi duality and modular invariance for the Dirac fermion on the torus arise when computing the entanglement of a single interval at finite temperature  Lokhande2015 ; Mukhi2018 ; Arias2020 .
@@ 2 Summary of results
$I_{3}^{(\alpha)}(x)=\frac{1}{\alpha-1}\log\left(\sum_{\vec{\varepsilon}\in\{0,\frac{1}{2}\}^{\alpha-1}}\frac{\Theta\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{0}\end{smallmatrix}\right]\left(\vec{0}|\frac{1}{\eta}\Omega\right)\Theta\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{0}\end{smallmatrix}\right]\left(\vec{0}|4\eta\Omega\right)}{\left|\Theta\left(\vec{0}|\Omega\right)\right|^{2}}\right)\,,$ (2.4)
where the subscript $\mathrm{f}$ is used to stress that we are considering fermionic entropies. Here $\eta$ is the universal critical exponent proportional to the square of the compactification radius of the underlying bosonic theory, which for $h=0$ reads $\eta=\frac{\arccos(-\Delta)}{\pi}$; $\Omega$ is the $(\alpha-1)\times(\alpha-1)$ period matrix of the Riemann surface $\mathcal{R}_{\alpha}$ with elements
@@
Note that it is customary to omit the characteristic when it is zero, i.e., $\Theta\equiv\Theta[\vec{0}]$. The behavior of the tripartite information for $\alpha=2$ is shown in Fig. 2. The curves for higher $\alpha=2$ are barely distinguishable from the ones for $\alpha=2$ so we do not present them.
For small $x$ we find the von Neumann tripartite information
@@
where we define the function Calabrese2011Entanglement
@@
We find, in particular, that the tripartite information is negative for $\frac{1}{4}\leq\eta\leq 1$ (i.e., $\Delta\geq-\frac{\sqrt{2}}{2}$ when $h=0$), while in the XXZ spin chain it is always non-negative. For the Rényi-$\alpha$ tripartite information in the case $\alpha=2,3$ we also find an unusual behavior: for certain values of $\Delta$, tripartite information changes the sign as a function of $x$ (see Fig. 3).
@@
(a) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(-0.3,0)$. The three subsystems are of equal length $|A|=|B|=|C|$, so that the cross ratio is $x=1/4$.
(a) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(-0.3,0)$. The three subsystems are of equal length $|A|=|B|=|C|$, so that the cross ratio is $x=1/4$.
(b) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(-0.3,0)$. The subsystem lengths are set to be $(|A|,|B|,|C|)=(2\ell,\ell,3\ell)$ for $\ell=1,2,3,\ldots$, so that the cross ratio is $x=1/2$. The lines R-NS and NS-R overlap.
@@
where $\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)$ is given by Eq. (2.10). For $\nu=2$ the tripartite information is equal to the one in the corresponding model obtained by Jordan-Wigner transformation, which is further equal to twice the one in the fermionic XXZ chain,
@@
where $I_{3}^{(\alpha;\mathrm{fXXZ})}$ is given by Eq. (2.4). For small $x$ and any $\nu$ we have computed the von Neumann tripatite information. The results are given in Sec. 5.2. The tripartite information can be positive, negative or zero, depending on model parameters.
@@ 3 Tripartite information
3.1 Fermions vs spins: spin structures—review
@@
for fermions. Note that (3.1) and (3.5) for spins do not differ only in the normalisation, but also in the space where they are defined, as we emphasized by using or not using bold fonts, and similarly for fermions. The entanglement entropies in terms of $\bm{\mathbf{R}}_{X}$ read
@@
which is instead an entangled state. The entropy of spins now equals $S_{\alpha}(A\cup C)=\log 2$. One could easily see that the relation (3.8) gives the correct entropy in both examples.
@@
Eq. (3.18) is exact on the lattice. Ref. Fagotti2010disjoint exploited this structure to compute the Rényi entropies in spin systems that are mapped to free fermions by the Jordan-Wigner transformation. Indeed, in that case, $\rho^{s,s^{\prime}}_{AC}$ turns out to be Gaussian for each pair $(s,s^{\prime})$, whereas the full reduced density matrix is not. When specialised to the ground state of critical chains, these notations make it manifest the correspondence pointed out in Ref. Coser2016Spin , according to which $(-1)^{4\vec{\varepsilon}\cdot\vec{\delta}}\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](A,B,C)$ is the partition function of the underlying fermionic quantum field theory on the Riemann surface associated with the Rényi-$\alpha$ entropy, with given boundary conditions along the cycles of the canonical homology basis. We now discuss this identification.
@@
$\mathrm{tr}\left[\prod\limits_{k=1}^{\alpha-1}\rho_{AC}^{s_{k}[\vec{\varepsilon}],s_{k}^{\prime}[\vec{\delta}]}\right]=Z_{\vec{\varepsilon},\vec{\delta}}=\int_{\vec{\varepsilon},\vec{\delta}}\mathcal{D}^{\dagger}\psi\mathcal{D}\psi\ e^{-S_{E}[\psi^{\dagger},\psi]},$ (3.22)
@@
3.2 Conformal critical systems—review
In this section we briefly review the behavior of the entanglement entropies in critical spin chains described by conformal field theory and the corresponding formalism. We focus on ground states in the thermodynamic limit. More details about the formalism for computing the entanglement entropies in QFT can be found in reviews Calabrese2009Entanglement ; Rangamani2017 .
@@ 3.2 Conformal critical systems—review
The relation between the tripartite information of $A,B,C$ and the mutual information $I_{2}^{(\alpha)}(A,C)=S_{\alpha}(A)+S_{\alpha}(C)-S_{\alpha}(AC)$ follows from Eqs (1.2) and (1.3)
@@ 3.3 Noncritical systems
$\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)=1\,.$ (3.30)
@@
$\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)=(-1)^{4\vec{\varepsilon}\cdot\vec{\delta}}$ (3.32)
@@
$H_{\mathrm{XY}}=\sum_{\ell}\left[(1+\gamma)\bm{\mathbf{\sigma}}_{\ell}^{x}\bm{\mathbf{\sigma}}_{\ell+1}^{x}+(1-\gamma)\bm{\mathbf{\sigma}}_{\ell}^{y}\bm{\mathbf{\sigma}}_{\ell+1}^{y}+2h\bm{\mathbf{\sigma}}_{\ell}^{z}\right]\;,$ (3.34)
@@ 4 Heisenberg XXZ model
4.1 Quantum field theory description—review
The XXZ spin chain (2.1) is critical in a range of parameters, e.g. in the absence of magnetic field $h$ the critical region is $\Delta\in[-1,1]$. Its low energy and large distance properties admit an effective Luttinger liquid description, developed in Refs Luther1975 ; Haldane1980 . For a pedagogical introduction to the effective quantum field theory description in terms of Thirring and free compactified boson Lagrangian we refer the reader to Ref. Affleck1988LesHouches .
The identification of the QFT is based on the Jordan-Wigner transformation. The identification starts with a perturbative treatment and is then lifted to a non-perturbative regime by fixing the QFT parameter phenomenologically, through the comparison of a QFT result and an exact lattice results for some quantity. It starts by identifying the non-interacting model with free Dirac fermions and treating the interacting part, quartic in Jordan-Wigner fermions, as a small perturbation. This yields the massless Thirring model, whose Lagrangian in the Euclidean space reads
@@ 4.1 Quantum field theory description—review
where the Dirac spinors have two components and the Dirac matrices satisfy $\{\gamma_{\mu},\gamma_{\nu}\}=\delta_{\mu,\nu}$ (e.g. $\gamma_{0}=\sigma^{y},\ \gamma_{1}=\sigma^{x}$). Bosonization is then used to map the Thirring model to the free boson compactified on a circle, given by the Lagrangian
@@
where the bosonic fields are periodic in the target space with period $2\pi$, i.e. $\varphi\equiv\varphi+2\pi$, and the compactification radius $R$ is related to the Thirring coupling constant through $R^{2}=1+\lambda$. Note that throughout the literature different conventions for the compactification radius are used. The one we use is associated to the T-duality $R\leftrightarrow 2/R$.
Next, symmetry arguments are used to conclude that there are no other relevant Lorentz-invariant interactions in a range of $R$, while the various non-renormalizible interactions can only lead to a renormalization of $R$. The compactification radius is then determined phenomenologically, by comparing the lattice result for some quantity, typically obtained with Bethe ansatz, with the field theoretical one. Various quantities can be used for this purpose, such as some critical exponents Luther1975 , finite-size spectrum Haldane1980 ; Alcaraz1987 ; Lukyanov1998 , compressibility Giamarchi2003book and susceptibility Eggert1994 . For zero magnetic field $h=0$ the identification is
@@
For a non-zero magnetic field the last equality in Eq. (4.3) does not hold, nor Eq. (4.4) for the spin correlation functions. In the presence of a magnetic field $\eta$ can be obtained by solving numerically integral equations obtained from Bethe ansatz Cabra1998 or through perturbative treatments Granet2019 .

### 11 Mar 2024 v3

@@ 1 Introduction
An important aspect that distinguishes the entropy of disjoint blocks from that of an interval is its sensitivity to duality mappings, in that such transformations generally preserve the locality of only a subset of operators (including the Hamiltonian density), breaking it instead into semilocality for the rest of them. This affects the meaning itself of subsystem, especially when it is not simply connected. A consequence of it on the behaviour of the entanglement entropies was firstly observed in the ground state of XY spin chains Igloi2010 ; Fagotti2010disjoint . At the critical point the tripartite information has a nontrivial scaling limit despite vanishing in the corresponding free-fermion model that emerges after a Jordan-Wigner transformation Casini2005 ; Casini2009reduced_density ; Casini2009Entanglement .
@@
Since the first step towards an analytical investigation into the entanglement entropies of spin blocks is often a mapping to quasiparticles, it is important to understand how such mappings affect the entropies and how to ultimately fix the lack of equivalence. In that respect, Ref. Fagotti2010disjoint proposed an expansion of the tripartite information based on the decomposition of the reduced density matrix into four pseudo reduced density matrices (pseudo-RDMs) that unfold the spin-flip symmetry. On the other hand, the critical spin-chain models studied in Ref. Fagotti2010disjoint admit a quantum field theory description of the low-energy and large-distance limit (the Ising model is described by a massless Majorana field theory, while the XX model by a Dirac one); as mentioned above, the Rényi entropy $S_{\alpha}$ of two disjoint intervals can then be written in terms of partition functions on particular Riemann surfaces. Ref. Coser2016Spin examined what the lattice procedure of Ref. Fagotti2010disjoint translates into in the quantum field theory description and established a one-to-one correspondence between the partition functions with given spin structures and the moments of the pseudo-RDMs introduced in Ref. Fagotti2010disjoint . It was argued, in particular, that the entropy of disjoint blocks of spins is obtained by summing over all spin structures of the fermion model on the underlying Riemann surface, i.e., over all combinations of antiperiodic, also called Neveu-Schwarz (NS), and periodic, also called Ramond (R), boundary conditions on the cycles of the surface. The entropy of Jordan-Wigner fermions, instead, corresponds to antiperiodic boundary conditions over all cycles (see also Ref. Camargo2021 ). As discussed in Ref. Headrick2010 in detail, indeed, the discrepancy between the entanglement entropy of two disjoint intervals for the free Dirac fermion Casini2005 ; Casini2009reduced_density ; Casini2009Entanglement and for the free boson compactified on a circle with the corresponding radius Calabrese2009Entanglement1 , in our notations $R=1$, can be traced back to the spin structure. The result for the boson matches, instead, that for the modular invariant Dirac fermion Calabrese2011Entanglement ; Coser2016Spin , which is obtained by summing over all boundary conditions. We mention that similar problems related to Bose-Fermi duality and modular invariance for the Dirac fermion on the torus arise when computing the entanglement of a single interval at finite temperature  Lokhande2015 ; Mukhi2018 ; Arias2020 .
@@ 2 Summary of results
$I_{3}^{(\alpha)}(x)=\frac{1}{\alpha-1}\log\left(\sum_{\vec{\varepsilon}\in\{0,\frac{1}{2}\}^{\alpha-1}}\frac{\Theta\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{0}\end{smallmatrix}\right]\left(\vec{0}|\frac{1}{\eta}\Omega\right)\Theta\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{0}\end{smallmatrix}\right]\left(\vec{0}|4\eta\Omega\right)}{\left|\Theta\left(\vec{0}|\Omega\right)\right|^{2}}\right)\,.$ (2.4)
Here $\eta$ is the universal critical exponent proportional to the square of the compactification radius of the underlying bosonic theory, which for $h=0$ reads $\eta=\frac{\arccos(-\Delta)}{\pi}$; $\Omega$ is the $(\alpha-1)\times(\alpha-1)$ period matrix of the Riemann surface $\mathcal{R}_{\alpha}$ with elements
@@
Note that it is customary to omit the characteristic when it is zero, i.e., $\Theta\equiv\Theta[\vec{0}]$. The behavior of the tripartite information for $\alpha=2$ is shown in Fig. 2. The curves for higher $\alpha$ are barely distinguishable from those for $\alpha=2$ so we do not present them.
We also obtain the von Neumann tripartite information in the limit of small $x$
@@
where $s_{\mathrm{vN}}$ is defined as Calabrese2011Entanglement
@@
We find, in particular, that the tripartite information is negative for $\frac{1}{4}\leq\eta\leq 1$ (i.e., $\Delta\geq-\frac{\sqrt{2}}{2}$ when $h=0$), while in the XXZ spin chain it is always non-negative. For the Rényi-$\alpha$ tripartite information with $\alpha=2,3$ we report an unusual behavior: for certain values of $\Delta$, tripartite information changes the sign as a function of $x$ (see Fig. 3).
@@
(a) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(-0.3,0)$. The three subsystems have equal lengths $|A|=|B|=|C|$, so the cross ratio is $x=1/4$. The dashed lines show the free case ($\Delta=0$) for comparison.
(a) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(-0.3,0)$. The three subsystems have equal lengths $|A|=|B|=|C|$, so the cross ratio is $x=1/4$. The dashed lines show the free case ($\Delta=0$) for comparison.
(b) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(-0.3,0)$. The subsystem lengths are set to be $(|A|,|B|,|C|)=(2\ell,\ell,3\ell)$ for $\ell=1,2,3,\ldots$, so the cross ratio is $x=1/2$. The lines R-NS and NS-R overlap. The dashed lines show the free case for comparison.
@@
where $\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)$ is given by Eq. (2.10). For $\nu=2$ the spin tripartite information matches the fermionic one, which, in turn, is equal to twice the tripartite information in the fermionic XXZ chain
@@
where $I_{3}^{(\alpha;\mathrm{fXXZ})}$ is given by Eq. (2.4). For small $x$ and any $\nu$ we have computed the von Neumann tripatite information. The results are given in Sec. 5.2. The tripartite information can be positive, negative or zero, depending on the model parameters.
@@ 3 Tripartite information
3.1 Fermions vs spins: spin structures
@@
for fermions. Note that (3.1) and (3.5) do not differ only in the normalisation, but also in the space where they are defined, as we emphasized by using or not using bold fonts, and similarly for (3.2) and (3.6). The entanglement entropies in terms of $\bm{\mathbf{R}}_{X}$ read
@@
which is instead an entangled state. The entropy of spins $S_{\alpha}(A\cup C)$ now equals $\log 2$. One could easily see that relation (3.8) gives the correct entropy in both examples.
@@
Eq. (3.18) is exact on the lattice. Ref. Fagotti2010disjoint exploited this structure to compute the Rényi entropies in spin systems that are mapped to free fermions by the Jordan-Wigner transformation. Indeed, in that case, $\rho^{s,s^{\prime}}_{AC}$ turns out to be Gaussian for each pair $(s,s^{\prime})$, whereas the full reduced density matrix is not. When specialised to the ground state of critical chains, these notations make it manifest the correspondence pointed out in Ref. Coser2016Spin , according to which $\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](A,B,C)$ is the partition function of the underlying fermionic quantum field theory on the Riemann surface associated with the Rényi-$\alpha$ entropy, with given boundary conditions along the cycles of the canonical homology basis. We now discuss this identification.
@@
$\mathrm{tr}\left[\prod\limits_{k=1}^{\alpha-1}\rho_{AC}^{s_{k}[\vec{\varepsilon}],s_{k}^{\prime}[\vec{\delta}]}\right]=Z_{\vec{\varepsilon},\vec{\delta}}=\int_{\vec{\varepsilon},\vec{\delta}}\mathcal{D}\psi^{\dagger}\mathcal{D}\psi\ e^{-S_{E}[\psi^{\dagger},\psi]},$ (3.22)
@@
3.2 Conformal critical systems
In this section we briefly review the behavior of the entanglement entropies in critical spin chains described by a conformal field theory, and the corresponding formalism. We focus on ground states in the thermodynamic limit. More details about the formalism for computing the entanglement entropies in QFT can be found in reviews Calabrese2009Entanglement ; Rangamani2017 .
@@ 3.2 Conformal critical systems
The relation between the tripartite information of $A,B,C$ and the mutual information $I_{2}^{(\alpha)}(A,C)=S_{\alpha}[A]+S_{\alpha}[C]-S_{\alpha}[A\cup C]$ follows from Eqs (1.2) and (1.3)
@@ 3.3 Noncritical systems
$\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right]=1\,.$ (3.30)
@@
$\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right]=(-1)^{4\vec{\varepsilon}\cdot\vec{\delta}}$ (3.32)
@@
$\bm{\mathbf{H}}_{\mathrm{XY}}=\sum_{\ell}\left[(1+\gamma)\bm{\mathbf{\sigma}}_{\ell}^{x}\bm{\mathbf{\sigma}}_{\ell+1}^{x}+(1-\gamma)\bm{\mathbf{\sigma}}_{\ell}^{y}\bm{\mathbf{\sigma}}_{\ell+1}^{y}+2h\bm{\mathbf{\sigma}}_{\ell}^{z}\right]\;,$ (3.34)
@@ 4 Heisenberg XXZ model
4.1 Quantum field theory description
The XXZ spin chain (2.1) is critical in a range of parameters, e.g., in the absence of magnetic field $h$ the critical region is $\Delta\in[-1,1]$. Its low energy and large distance properties admit an effective Luttinger liquid description, developed in Refs Luther1975 ; Haldane1980 . For a pedagogical introduction to the effective quantum field theory description in terms of Thirring and free compactified boson Lagrangian we refer the reader to Ref. Affleck1988LesHouches .
The identification of the QFT is based on the Jordan-Wigner transformation. It is first established perturbatively and is then lifted to a non-perturbative level by fixing the QFT parameters phenomenologically, through the comparison bewteen QFT predictions and exact lattice results. Specifically, the non-interacting model is identified with a free Dirac fermion and the interacting part, quartic in the Jordan-Wigner fermions, is treated as a small perturbation. This yields the massless Thirring model, whose Lagrangian in the Euclidean space reads
@@ 4.1 Quantum field theory description
where the spinors $\psi$ have two components and the Dirac matrices satisfy $\{\gamma_{\mu},\gamma_{\nu}\}=\delta_{\mu,\nu}$ (e.g. $\gamma_{0}=\sigma^{y},\ \gamma_{1}=\sigma^{x}$). Bosonization is then used to map the Thirring model to the free boson compactified on a circle, given by the Lagrangian
@@
where the bosonic fields are periodic in the target space with period $2\pi$, i.e. $\varphi\equiv\varphi+2\pi$, and the compactification radius $R$ is related to the Thirring coupling constant through $R^{2}=1+\lambda$. Note that throughout the literature different conventions for the compactification radius are used. In the one we use the T-duality reads $R\leftrightarrow 2/R$.
Next, symmetry arguments are used to conclude that there are no other relevant Lorentz-invariant interactions in a range of $R$, while the various non-renormalizible interactions can only lead to a renormalization of $R$. The compactification radius is finally determined phenomenologically, by comparing the lattice result for some quantity, typically obtained within Bethe ansatz, with the field theoretical one. Various quantities can be used to this purpose, such as some critical exponents Luther1975 , finite-size spectrum Haldane1980 ; Alcaraz1987 ; Lukyanov1998 , compressibility Giamarchi2003book and susceptibility Eggert1994 . For zero magnetic field $h=0$ the identification is
@@
For a non-zero magnetic field the last equality in Eq. (4.3) does not hold, nor Eq. (4.4) for the spin correlation functions. In that case $\eta$ is obtained numerically by solving integral equations Cabra1998 or through perturbative methods Granet2019 .